OpenStudy (stamp):

Simplify \[\sqrt{6u^2y}\sqrt{2u^9y^2}\]

OpenStudy (stamp):

Oh god what is the mathtext for squar root

OpenStudy (vishweshshrimali5):

\sqrt

OpenStudy (vishweshshrimali5):

Sorry: \sqrt{}

OpenStudy (anonymous):

\[\sqrt{6u^2y}\sqrt{2u^9y^2}\]

OpenStudy (vishweshshrimali5):

^^^ Like this ;)

OpenStudy (stamp):

\[\sqrt{6u^2y}\sqrt{2u^9y^2}\]

OpenStudy (vishweshshrimali5):

Yeah

OpenStudy (stamp):

I got my answer, would someone please give me a verification?

OpenStudy (vishweshshrimali5):

Sure

OpenStudy (stamp):

\[u^5y\sqrt{12uy}=2u^5y\sqrt{3uy}\]

OpenStudy (vishweshshrimali5):

Correct :)

OpenStudy (vishweshshrimali5):

Good work !